calculate the energy of a photon of radiation with a wavelength of 6.4 * 10^-7
1
Expert's answer
2020-10-19T14:27:11-0400

According to the equation:

E = hc / λ

where E - energy, h - Planck constant (6.63 × 10-34 J s), c - speed of light (3 × 108 m/s), λ - photon wavelength.

From here:

E = 6.63 × 10-34 J s × 3 × 108 m/s / 6.4 × 10-7 m = 3.1 × 10-19 J


Answer: 3.1 × 10-19 J